a closed vessel with rigid walls contains 1 mol of uranium and 1 mol of air at 298k.considering complete decay of uranium to lead,the ratio of the final pressure o initial pressure of system at 298k is

Grade:12th pass

1 Answers

Arun
5 years ago
The number of gaseous moles initially = 1 mol....... The number of gaseous moles finally = 1 + 8 mol.....so the ratio = 9/1